4. Two bulbs, 1.00 L and 1.50 L in volume, connected by a stopcock, are filled, respectively,
with argon at 0.75 atm and helium at 1.20 atm at the same temperature. Calculate the mole
fraction of each gas, total pressure and partial pressure of each gas after the stopcock is
opened. Assume ideal gas behaviour.
1
Expert's answer
2020-12-14T14:50:07-0500

PIVI=P2V2

Volume of Argon= 1.00L and its pressure=0.75atm

Volume of Helium=1.50L and its pressure=1.20atm

Total volume=(1.00+1.50)L=2.5L


P2 Argon= 1×0.75/2.5=0.3atm (partial pressure of Argon gas)

P2 Helium=1.5×1.20/2.5=0.72atm (partial pressure of Helium gas)

Total pressure=(0.3+0.72)atm

=1.02atm

P2 Argon/P total= n Argon/n Total

0.3/1.02=0.2941 mole fraction of Argon

P2 Helium/P total= n Helium/n Total

0.72/1.02=0.7059 mole fraction of of Helium